OpenStudy (anonymous):

You are playing a game where you have 1 - 4 a chance of winning $500, 1 - 4 chance of winning $200 and 1 - 2 chance of losing $375. What is your expected payoff?

OpenStudy (paxpolaris):

\[\Large \frac14 \cdot 500 +\frac 14 \cdot 200 \ -\frac 12 \cdot 375\]

OpenStudy (anonymous):

A) -$12.50 B) $108.33 C) $325 D) $362.50 Hint

OpenStudy (paxpolaris):

do the math...

OpenStudy (anonymous):

It's -$12.50

OpenStudy (paxpolaris):

\[\Huge \checkmark\]